OpenStudy (anonymous):

use properties of exponents to simplify the expression. write with positive exponent (8x^3)^1/3/2x^4

OpenStudy (anonymous):

(x^a)^b = x^ab x^a/x^b = x^(a-b)

OpenStudy (anonymous):

Is this: \[((8x^3)^1) / (3/(2x^4))\]

OpenStudy (anonymous):